OpenStudy (anonymous):

Is 8 a solution of x < -1 or x > 7?

OpenStudy (anonymous):

yes or no?

OpenStudy (tkhunny):

Is 8 less than -1? Is 8 greater than 7? Did you answer "yes" to either of those?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

There you have it! What's the next problem?

OpenStudy (anonymous):

Which of the following is equivalent to |4x + 5| = 88?

OpenStudy (anonymous):

-4x + 5 = 88 4x + 5 = -88 4x + 5 = 88 4x + 5 = 88 or 4x + 5 = -88

OpenStudy (anonymous):

c?

OpenStudy (tkhunny):

a, b, and c aren't even candidates. |88| = 88 |-88| = 88 Unless it's zero over there on the right-hand side, there MUST be TWO parts of the solution.

OpenStudy (anonymous):

i see Thanks a lot!!